OpenStudy (anonymous):

solve the proportion t/16=40/50 A.t=26 B.t=28 C.t=1.28 D. t=12.8 @OrangeMaster

OpenStudy (anonymous):

You have to cross multiply.

OpenStudy (anonymous):

t 45 ___ = ____ 40 50

OpenStudy (anonymous):

how do i do it??

OpenStudy (anonymous):

50t=1800

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Divide both sides by 50 to get t=36. I don't see that as an answer choice? I'm confused. Just wait a sec.

OpenStudy (anonymous):

@CGGURUMANJUNATH, can you help me. I thought that I did it right.

OpenStudy (cggurumanjunath):

:)

OpenStudy (cggurumanjunath):

u got it . @OrangeMaster

OpenStudy (anonymous):

@keke_luvs_u, I don't see t=36 as an answer choice.

OpenStudy (cggurumanjunath):

a should be 36 ,not 26 @keke_luvs_u

OpenStudy (anonymous):

i checked my asnwers and 36 isn't a choice

OpenStudy (anonymous):

Yeah, maybe that's the problem. Well 36 is correct.

OpenStudy (cggurumanjunath):

then check the q again ! @keke_luvs_u

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

I'm standing by 36. I know that it is the correct answer.

OpenStudy (anonymous):

Where did you get the question?

OpenStudy (cggurumanjunath):

I'm standing by 36. I know that it is the correct answer. me too.

OpenStudy (anonymous):

connexous

OpenStudy (anonymous):

Then webmail your teacher, and see if the question was a mistake.

OpenStudy (anonymous):

the given is t/16=40/50 not t 45 ___ = ____ 40 50

OpenStudy (cggurumanjunath):

good to go @mz_samosino

OpenStudy (cggurumanjunath):

yes @OrangeMaster ;u had copied the q wrongly.

OpenStudy (anonymous):

It doesn't matter. Multiplication is the same no matter how you arrange the numbers. My answer is still correct.

OpenStudy (anonymous):

....

OpenStudy (cggurumanjunath):

but ur method for ur problem is right @OrangeMaster

OpenStudy (cggurumanjunath):

ur lhs is right @OrangeMaster

OpenStudy (cggurumanjunath):

but rhs should have been 16/50 and not 45/50

OpenStudy (anonymous):

@keke_luvs_u dont you understand cross multipliation

OpenStudy (cggurumanjunath):

lhs-left hand side . rhs-right hand side .

OpenStudy (anonymous):

I think we can all agree that mz_samosino had the right question, but I had the correct answer as well.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

Either way you look at it, there must be something wrong on your end keke_luvs_u.

OpenStudy (cggurumanjunath):

t =12.8

OpenStudy (anonymous):

OK I see where I messed up.

OpenStudy (anonymous):

Everybody should give a medal to @mz_samosino. She did really good.

OpenStudy (anonymous):

so the answer is 36 or not?

OpenStudy (anonymous):

ohh so 12.8!

OpenStudy (anonymous):

no

OpenStudy (anonymous):

not 36

OpenStudy (anonymous):

I messed up when coping the question.

OpenStudy (anonymous):

@keke_luvs_u study the cross multiplication you are troubling there i already give you that picture earlier but still you didn't undertand